Circular No. 6858

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ab IN NGC 4704 J.-y. Wei, on behalf of the Beijing Astronomical Observatory supernova

survey (via W.-d. Li, Berkeley), reports the discovery of an apparent supernova (mag 16.1) on Apr. 1.67 UT. SN 1998ab is located at α = 12^{h}48^{m}47^{s}.24, δ = +41^{o}55′28′′.3 (equinox 2000.0), which is 2′′ east and 12′′ north of the center of NGC 4704. SN 1998ab is not visible on a CCD image taken on Mar. 14.

P. Garnavich, S. Jha, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, write: “A spectrum was taken of SN 1998ab by P. Berlind and M. Calkins with the 1.5-m Tillinghast telescope on Apr. 4.4 UT. The spectrum shows it to be an unusual type-Ia supernova, similar to SN 1991T before maximum. Deep absorption features observed at 437 and 507 nm are the strongest features on a blue continuum and correspond to Fe III. No Si II is clearly visible. A broad absorption feature at 381 nm is likely to be Ca II. Narrow emission lines from the host galaxy give a redshift of 8250 km/s. Concurrent images taken by E. Barton with the Whipple Observatory’s 1.2-m telescope yield estimated magnitudes B = 16.6 and V = 16.4. The supernova is offset 9′′.2 east and 12′′.1 north from the nucleus of NGC 4704. Spectroscopic and photometric observations of this unusual supernova are encouraged.”

AX J0051–722 J. C. Lochner, Goddard Space Flight Center and Universities Space

Research Association, reports: “After fading steadily from its appearance in late Nov. 1997, the transient x-ray system AX J0051–722 (IAUC 6803) was observed to have rebrightened in an observation by RXTE on 1998 Mar. 25. Weak pulsations at 59 s from XTE J0055–724 = 1SAX J0054.9– 7226 (IAUC 6818) are also evident in these data. The estimated 2–10-keV flux of AX J0051–722 for this observation is 5.7 × 10^{−11} erg s^{−1} cm^{−2}. An observation by RXTE on Mar. 31 showed very strong pulsations at 91.0 ± 0.1 s, and a decrease in the flux to 4.0 × 10^{−11} erg s^{−1} cm^{−2}. The spectrum can be described by a cutoff power law with photon index of 0.80 and cutoff energy of 14.0 keV. This reappearance suggests a possible orbital period of this Be/neutron-star system of ∼ 110 days. Continued optical monitoring of the system is strongly urged.”

1998 April 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green